# Follow-Bench: A Unified Motion Planning Benchmark for Socially-Aware Robot Person Following

### v2

@@ VI-B Detailed Experimental Results
Here, we detail our experimental results of every single scenario under different human numbers, occupied conditions and following distances. Fig. 11 and Fig. 12 show the performance under varying human numbers in dynamic crowd flows and environmental layouts, respectively. Fig. 13 shows the performance under different environmental occupied conditions in environmental layouts. Fig. 14 and Fig. 15 show the performance under varying following distances in dynamic crowd flows and environmental layouts, respectively. More evaluation results of other evaluation metrics can be found in the released code.

### v3

@@ VI-A Baselines and Simulation Settings
In all simulated scenarios, the baseline configuration models each human as an omni-directional dynamic cylinder (radius: , max speed: ) and the robot as a differential-drive platform () equipped with a 2D LiDAR ( range), subject to translational velocity and acceleration limits. As for varied human densities and environmental occupancy conditions, the maximum pedestrian densities, obstacle numbers, and minimum passage widths are determined empirically by observing conditions under which simulated pedestrians become prone to congestion.
@@ VI-B Detailed Experimental Results
Here, we detail our experimental results of every single scenario under different numbers of humans, occupied conditions and following distances. Fig. 11 and Fig. 12 show the performance under varying numbers of humans in dynamic crowd flows and environmental layouts, respectively. Fig. 13 shows the performance under different environmental occupied conditions in environmental layouts. Fig. 14 and Fig. 15 show the performance under varying following distances in dynamic crowd flows and environmental layouts, respectively. More evaluation results of other evaluation metrics can be found in the released code.